OpenStudy (anonymous):

factor p 2 + 18p + 32

TheSmartOne (thesmartone):

Hint: 16 x 2 = 32 16 + 2 = 8

OpenStudy (alexandervonhumboldt2):

you are less than a first grader @TheSmartOne well joke of course but 16+2=18 , not 8.

TheSmartOne (thesmartone):

It was a typo. I meant 16 + 2 = 18

TheSmartOne (thesmartone):

With those hints the asker should get the answer. If he doesn't, he better ask for more help.